Feed A has 10% protein and feed B has 4% protein. How many pounds of each are mixed to get 85 pounds of feed with 8% protein. Write and solve the system.
Answer by richwmiller(17219)

a+b=85,
.10a + .04b=.08*85
solve by substitution
b=85-a
.10a+.04(85-a)=.08*85
.10a+.04*85-.04a=.08*85
.06a=.04*85
a=(.04/.06)*85
a=2/3*85
a=56.6667, b=28.3333